Calculate the fundamental stretching frequency and wavenumber for a C–H bond, considering that the atoms vibrate independently of other groups on a carbon atom and that the force constant is 4800/nm



1
Expert's answer
2021-01-18T01:44:26-0500

The fundamental stretching wavenumber "\bar\nu" can be calculated from the force constant "k" (4800 N/m, or kg/s2) using the following equation:


"\bar\nu = \frac{1}{2\pi c}\sqrt{\frac{k}{\mu}}" ,

where "c" is the speed of light 3·1010 cm/s, "\mu" is the reduced mass and "k" is the force constant.


The reduced mass of C and H is (taken from the molar masses in kg/mol and Avogadro's number 6.022·1023 mol-1):

"\mu= \frac{m_Cm_H}{m_C + m_H} = \frac{12·1·10^{-3}}{(12 +1)·6.022·10^{23}} = 1.53·10^{-27}" kg.

Finally, the fundamental stretching wavenumber is:

"\bar\nu = \frac{1}{2\pi·3·10^{10}}\sqrt\frac{4800}{1.53·10^{-27}} = 9397" cm-1.

And the frequency:

"\nu = \bar{\nu}c = 2.819·10^{14}" s-1.